OpenStudy (anonymous):

simplify radical8 divied by radical 100

OpenStudy (lgbasallote):

\[\huge \frac{\sqrt 8}{\sqrt{100}} \implies \sqrt{\frac 8{100}}\] does that help?

OpenStudy (anonymous):

i know that part

OpenStudy (lgbasallote):

do you know how to simplify 8/100 into lowest terms?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

\[\huge \frac 8{100} \implies \frac{8 \div 4}{100 \div 4} \implies \frac 2{25}\] got it?

OpenStudy (anonymous):

yes thank you

OpenStudy (lgbasallote):

good.. so \[\huge \sqrt{\frac 8{100}} \implies \sqrt{\frac 2{25}}\] agree?